# Twinning Complex Networked Systems: Data-Driven Calibration of the mABCD Synthetic Graph Generator

### v1

@@ 1 Introduction
To accurately characterise dynamics in complex networked systems, it is necessary to adopt an approach that differentiates between the various types of relationships present. A natural way to achieve this is through the use of multilayer networks [15], a system defined as a collection of $\ell$ layers, where each layer represents a distinct mode of interaction encoded by its corresponding graph. The resulting dynamics are governed not only by intra-layer edges, describing interactions within a given relationship type, but also by inter-layer couplings, which connect the same entity across different layers. This structure enables modelling of complex spreading mechanisms that are obscured in monoplex representations [6].
@@ 2 Preliminaries
In order to address the problem studied in this paper, we first introduce the fundamental concepts used throughout. In particular, we briefly review multilayer networks, outline the general concept for the configuration inference task, and describe the synthetic graph generator employed.
@@ 2.1 Multilayer networks
As stated in the introduction, heterogeneity of complex networked systems is a crucial property which cannot be neglected. Therefore, the problem tackled in this work aims to take into account this property by utilising multilayer networks. Following [15], we formalise the concept of a multilayer network below. Nevertheless, before we do this, we clarify that for a given $\in{\mathbb{N}}=\{1,2,\ldots\}, [n][n$] is used to denote the set consisting of the first $n$ natural numbers, that is, $n]=\{1,2,\ldots,n\}$.
@@ 2.2 Complex Networked System Analysis with Twinning Approach
Utilising the example already introduced in Sec. 1, a conceptual framework guiding the problem is presented here. Suppose one is to evaluate several “what-if” scenarios that cover interventions on a social network to design effective strategies to control misinformation spread, given an initial dataset of ground truth interactions on the platform.
@@ 2.3 The Synthetic Network Generators — mABCD
Synthetic graph generators with explicit community structure play a central role in the development and evaluation of analysis methods for complex networked systems, particularly in settings where large-scale empirical datasets with reliably annotated ground-truth communities are scarce. To address this limitation, several benchmark models have been proposed to generate artificial networks that resemble real-world systems while retaining full control over their structural properties. A prominent example is the LFR benchmark model [17, 18], which allows for heterogeneous degree distributions and community sizes and has become a standard tool for evaluating community detection algorithms.
More recently, the Artificial Benchmark for Community Detection (ABCD) model [12] has been introduced as a scalable [10] and theoretically tractable alternative to LFR. Undirected variants of ABCD generate networks with comparable structural properties, while offering improved computational performance and greater flexibility in interpolating between well-defined community structure and random graphs [10]. Owing to these properties, ABCD has been extensively analysed from a theoretical perspective, including studies of its modularity behaviour [11] and self-similarity properties [3]. The modular design of the model has further enabled a range of extensions, including outliers [13], overlapping communities [2], hypergraphs [14], and, most relevant to this work, multilayer networks [16]. The multilayer extension, mABCD, generalises the ABCD framework to systems in which multiple interaction layers jointly shape network dynamics. Rather than aiming to exactly reproduce a specific empirical network, mABCD is intended to generate ensembles of multilayer graphs that preserve key structural characteristics of a target system, and can therefore be interpreted as coarse-grained digital twins of complex networked systems.
In this work, we adopt mABCD as the synthetic generator underlying our experimental analysis. Therefore, to focus on the aspects most relevant to the configuration inference problem considered here, a brief review of its core principles is necessary. In principle, the mABCD  network generation comprises of six subsequent phases. The process is governed by global and local (independent for each layer) parameters (Tab. 1). which affect various properties of the model, see Sec. 3 for details.
Table 1: Global and local parameters of mABCD.
@@
$d$ $\mathbb{N}$} Dimension of reference layer
@@
Local parameters (independent for each layer)
@@
$\tau_{i}$ [1,1][-1,1] Correlation coefficient between degrees and labels
@@
$\delta_{i}$ $\mathbb{N}$} Min. degree at least $\delta_{i}$
$\Delta_{i}$ ${\mathbb{N}}\;(1\leq\delta_{i}\leq\Delta_{i}<n)$ Max. degree at most $\Delta_{i}$
@@
$s_{i}$ $\mathbb{N}$} Min community size at least $s_{i}$
$S_{i}$ ${\mathbb{N}}\;(\delta<s_{i}\leq S_{i}\leq n)$ Max community size at most $S_{i}$
@@
The first five steps are carried out independently for each of the $\ell$ layers, while the sixth one binds the generated graphs into a single multilayer network. In the first phase, the so-called active nodes (i.e., those that will not be isolated) are determined. Each active node is endowed its degree in the second step. During the third step, communities are created and populated with nodes based on the latent layer, which is implicitly generated by the routine to serve as a proxy for actors’ features that typically shape connections in complex networked systems. Phase four focuses on connecting nodes within other members of the same community as well as establishing inter-community (i.e., background) connections. These graphs are subsequently simplified to remove possible self-loops and multiedges while preserving the intended structural properties. The final phase performs a series of edge rewirings to achieve the targeted edge correlations between each pair of layers. For a detailed description, please see [16].
@@ 3 Proposed mABCD Configuration Retrieval Method
Suppose that we are given a multilayered network following Def. 1. Our goal is to generate a digital twin, a multilayered network $\hat{G}$ generated by mABCD with associated graphs $\hat{G}_{i}=(\hat{V}_{i},\hat{E}_{i})$ for $i\in[\ell]$, that mimics the original network $G$ as best as possible. In order to do it, one needs to appropriately select the parameters of the mABCD model and then measure the “success”, i.e., how close $G$ to $\hat{G}$ is. We will measure the success by computing a divergence score $\mathcal{D}(\hat{G},G)$; the smaller the score, the better the fit is obtained.
@@
3.0.1 Dimension of reference layer ($d$).
Communities in mABCD are independently generated in each layer, but the desired correlations are obtained via a hidden, low-dimensional geometric reference layer. Nodes close to each other in the reference layer end up in the same community with a larger probability. In particular, the dimension of a reference layer $d$ affects many properties of the generated network. However, it seems not so easy to guess the right value so we leave it as a parameter that eventually can be tuned for the best outcome, that is, the smallest divergence score $\mathcal{D}(\hat{G},G)$. Understanding the influence of this hyperparameter and selecting the optimal value is left as an open problem (Open Problem 2).
This geometric approach to community structure is justified by the fact that latent geometric spaces are widely believed to shape complex networks (e.g., social media networks shaped by users’ opinions, education, knowledge, interests, etc.). These latent spaces have been successfully employed for many years to model and explain network properties such as self-similarity [20], homophily, and aversion [9]. For more details, we direct the reader to the survey [4] or the book [19].
@@
To measure correlations between edges in different layers, we define $\mathbf{R}$, a $\ell\times\ell$ matrix in which elements $r_{i,j}\in[0,1]$ ($i,j\in[\ell]$) capture correlation between edges present in layers $i$ and $j$. For any $i,j\in[\ell]$, let $E_{i}^{j}$ be the set of edges that are present in layer $i$, involving actors that are also active in layer $j$. Entries $r_{i,j}$ in $\mathbf{R}$ are computed using the following formula:
$r_{i,j}=\frac{|E_{i}^{j}\cap E_{j}^{i}|}{\min\{|E_{i}^{j}|,|E_{j}^{i}|\}}.$
If $\min\{|E_{i}^{j}|,|E_{j}^{i}|\}=0$, then we leave $r_{i,j}$ undefined.
Note that $r_{i,i}=1$ for any $i\in[\ell]$ and $r_{i,j}=r_{j,i}$ for $1\leq i<j\leq\ell$. The maximum value of $1$ is attained when edges in one of the layers form a subset of edges in the other layer. The minimum value of 0 is attained when the two sets of edges in the corresponding layers are completely disjoint.
Extracting $A=\mathbf{R}$ from the original network $G$ is straightforward. The mABCD model aims to produce a synthetic network with the desired edge correlation matrix $A$, but this is not guaranteed, and some small error always occurs. Let $B$ be the corresponding edge correlation matrix for the obtained digital twin $\hat{G}$. The divergence score is simply the normalised to $[0,1]$ Frobenius norm between the two matrices:
$\mathcal{D}_{\mathbf{R}}(\hat{G},G)=\frac{\lVert A-B\rVert_{F}}{\sqrt{\ell(\ell-1)}}=\sqrt{\frac{1}{\ell(\ell-1)}\sum_{i,j\in[\ell]}(a_{ij}-b_{ij})^{2}}=\sqrt{\frac{1}{\binom{\ell}{2}}\sum_{1\leq i<j\leq\ell}(a_{ij}-b_{ij})^{2}}.$
@@
To estimate $\tau_{i}$, which relies on the ordering of actor labels, we first sort nodes with respect to their total degree (sum of degrees over all layers), breaking ties randomly if needed; that is, the node with the highest total degree is assigned label $n$, the highest possible label. Then, for any $i\in[\ell]$, we compute the Kendall rank correlation coefficient $\tau_{i}$ between these labels and the corresponding degree sequence in layer $i$. Moreover, only nodes with a positive degree are taken into account to obtain the correlation value in order to reduce the noise. In particular, inactive nodes are ignored as they should be.
To measure the quality of the fit, one can compute the matrix $A=(a_{ij})$, where $a_{ij}\in[-1,1]$ is the Kendall rank correlation coefficient between layers $i$ and $j$ for the real network $G$. Clearly, only nodes that are active in both layers (i.e., nodes in $V_{i}\cap V_{j}$) are considered. Once a digital twin is generated, matrix $B$ is computed analogously, but this time for the twin $\hat{G}$. As before, the divergence score is simply the (normalised) Frobenius norm between the two matrices:
$\mathcal{D}_{\tau}(\hat{G},G)=\frac{\lVert A-B\rVert_{F}}{\sqrt{4\ell(\ell-1)}}=\sqrt{\frac{1}{4\binom{\ell}{2}}\sum_{1\leq i<j\leq\ell}(a_{ij}-b_{ij})^{2}}.$
@@
This family of parameters models correlations between communities formed by nodes that are active in two different layers. These parameters seem to be the most challenging to retrieve, as we do not have access to the latent biscuit-like reference layer used during network generation. However, once a digital twin is generated, one can easily measure the success by computing the corresponding divergence score $\mathcal{D}_{r}(\hat{G},G)$. The relationship between the parameters $r_{i}$ and the divergence score is not clear, but one can minimise $\mathcal{D}_{r}(\hat{G},G)$ by efficiently searching the parameter space.
The proposed method for predicting $r$ bases on the access to the previously estimated remaining part of the mABCD configuration and the $A$ matrix of the original network. The objective function creates a given number of candidate twins according to the evaluated $r$ supplemented by the already known rest of the parameters. Next, it computes a $B$ matrix for each of them and returns the mean divergence score $\mathcal{D}_{r}$ over them. This value is then minimised during the Bayesian optimisation process. By sequentially evaluating the objective function at values of $r$ selected based on the outcomes of previous evaluations, the method seeks to identify the best-matching value of $r$. The implementation employed, provided by the skopt library, accounts for the non-deterministic behaviour of mABCD by incorporating Gaussian noise into the surrogate model. A typical optimisation trajectory for fitting $r$ is shown in Fig. 2.
Figure 2: A typical optimisation trajectory projected on $\mathfrak{R}^{2}$ space of $r$ obtained during the experimental part. Note the trade-off between exploration and exploitation that the employed method preserves, by which it avoids getting stuck in local minima.
Let us then concentrate on the divergence score. For each layer $i\in[\ell]$ in the real network $G$, we independently run some stable clustering algorithm (in experiments, the Greedy Modularity Optimisation [5]) to get a partition $\mathcal{P}_{i}$ that identifies its community structure. Then, one can compute the matrix $A=(a_{ij})$, where $a_{ij}\in[0,1]$ is the adjusted mutual information (AMI) between partitions $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ induced by nodes that are active in both layers, that is, partitions induced by $V_{i}\cap V_{j}$. Matrix $B$ is computed analogously but for the twin $\hat{G}$ and the divergence score is defined as follows:
$\mathcal{D}_{r}(\hat{G},G)=\frac{\lVert A-B\rVert_{F}}{\sqrt{\ell(\ell-1)}}=\sqrt{\frac{1}{\binom{\ell}{2}}\sum_{1\leq i<j\leq\ell}(a_{ij}-b_{ij})^{2}}.$
@@
The mABCD model generates, by design, sequences following a power-law with parameters provided as the input. Hence, the only discrepancy between the original network $G$ and its synthetic digital twin $\hat{G}$ comes from fitting the power-law distribution into a given real network degree sequence. To measure the quality of the fit, we might use the Kolmogorov–Smirnov test. This test is a nonparametric statistical test used to determine if a sample comes from a specific distribution (one-sample) or if two samples come from the same distribution (two-sample) by measuring the maximum difference between their cumulative distribution functions (CDFs). In our scenario, it implies the following divergence score of the distribution from the real network from the theoretical distribution:
@@
For this short paper, for simplicity, we assume that the degree sequence of a real multilayered network $G$ is close to a power law. But not all networks have degree sequences of this nature. To solve this issue, the mABCD model is flexible and allows the degree sequences to be directly injected into the model. Since mABCD is using the classical configuration model to generate edges in each layer, with this more advanced option, the degree sequence of the digital twin $\hat{G}$ would match exactly the original network $G$. We leave an implementation of this extension into our framework and an investigation of its consequences on the quality of the digital twin for the future (Open Problem 3).

### v2

@@ 1 Introduction
To accurately characterise dynamics in complex networked systems, it is necessary to adopt an approach that differentiates between the various types of relationships present. A natural way to achieve this is through the use of multilayer networks [16], a system defined as a collection of $\ell$ layers, where each layer represents a distinct mode of interaction encoded by its corresponding graph. The resulting dynamics are governed not only by intra-layer edges, describing interactions within a given relationship type, but also by the presence of the system’s actors across different layers, through which they serve as inter-layer bridges. This structure enables modelling of complex spreading mechanisms that are obscured in monoplex representations [7].
@@ 2 Preliminaries
To address the problem studied in this paper, we first introduce the fundamental concepts. We briefly review multilayer networks, outline the general concept for the configuration inference task, and describe the synthetic graph generator employed.
@@ 2.1 Multilayer networks
As stated in the introduction, heterogeneity of complex networked systems is a crucial property which cannot be neglected. Therefore, the problem tackled in this work aims to take into account this property by utilising multilayer networks. Following [16], we formalise the concept of a multilayer network below. Nevertheless, before we do this, we clarify that for a given $\in{\mathbb{N}}=\{1,2,\ldots\}, [n][n$] is used to denote the set consisting of the first $n$ natural numbers, that is, $n]=\{1,2,\ldots,n\}$. Furthermore, for a matrix $\mathbf{X}$, we denote its $(i,j)$-th entry by $\mathbf{X}_{ij}$.
@@ 2.2 Complex Networked System Analysis with Twinning Approach
Utilising the example introduced in Sec. 1, a conceptual framework guiding the problem is presented here. Suppose one is to evaluate several “what-if” scenarios that cover interventions in a social network to design effective strategies to control the spread of misinformation, given an initial dataset of ground-truth interactions on the platform.
@@ 2.3 The Synthetic Network Generators — mABCD
Synthetic graph generators with explicit community structure play a central role in the development and evaluation of analysis methods for complex networks, particularly in settings where large-scale empirical datasets with reliably annotated communities are scarce. To address this limitation, several benchmark models have been proposed to generate artificial networks that resemble real-world systems while retaining full control over their structural properties. A prominent example is the LFR model [18] and its extensions [5], which allows for heterogeneous degree distributions and community sizes and has become a standard tool for evaluating community detection algorithms.
More recently, the Artificial Benchmark for Community Detection (ABCD) model [13] has been introduced as a scalable [11] and theoretically tractable alternative to LFR. Undirected variants of ABCD generate networks with comparable structural properties, while offering improved computational performance and greater flexibility in interpolating between well-defined community structure and random graphs [11]. Owing to these properties, ABCD has been extensively analysed from a theoretical perspective, including studies of its modularity behaviour [12] and self-similarity properties [3]. The modular design of the model has further enabled a range of extensions, including outliers [14], overlapping communities [2], hypergraphs [15], and multilayer networks [17]. The multilayer extension, mABCD, generalises the ABCD framework to systems in which multiple interaction layers jointly shape network dynamics. Rather than aiming to exactly reproduce a specific empirical network, mABCD is intended to generate ensembles of graphs that preserve key structural characteristics of a target system, and can therefore be interpreted as coarse-grained digital twins of complex networked systems.
In this work, we adopt mABCD as the synthetic generator underlying our experimental analysis. Therefore, to focus on the aspects most relevant to the configuration inference problem considered here, a brief review of its core principles is necessary. In principle, the mABCD network generation comprises of six subsequent phases. The process is governed by global and local (independent for each layer) parameters presented briefly in Tab. 1. For a detailed description, please refer to [17].
Table 1: Global and local parameters of mABCD denoted together as $\Theta$.
@@
$d$ $\mathbb{N}$} Dimension of the biscuit-like reference layer
@@
Local parameters (independent for each $i$ layer)
@@
$\tau_{i}$ [1,1][-1,1] Correlation coef. between degrees and actors’ labels
@@
$\delta_{i}$ $\mathbb{N}$} Min. degree
$\Delta_{i}$ ${\mathbb{N}}\;(1\leq\delta_{i}\leq\Delta_{i}<n)$ Max. degree
@@
$s_{i}$ $\mathbb{N}$} Min. community size
$S_{i}$ ${\mathbb{N}}\;(\delta<s_{i}\leq S_{i}\leq n)$ Max. community size
@@
The first five steps are carried out independently for each of the $\ell$ layers, while the sixth one binds the generated graphs into a single multilayer network. In the first phase, the so-called active nodes (i.e., those that will not be isolated) are determined. Each active node is endowed with its degree in the second step. During the third step, communities (understood as in [6]) are created and populated with nodes. This is done with the help of the reference latent layer, which is implicitly generated during the network generation process and discarded afterwards. It serves as a proxy for actors’ features, which typically shape connections in complex networked systems. The layer can be visualised as a $d$-dimensional biscuit, with raisins representing actors: the closer two raisins are, the more likely the corresponding actors are to be assigned to the same community. Phase four focuses on connecting nodes within the communities as well as establishing inter-community (i.e., background) links. These graphs are subsequently simplified to remove possible self-loops and multiedges while preserving the intended structural properties. The final phase performs a series of edge rewirings to achieve the targeted edge correlations between each pair of layers.
@@ 3 Proposed mABCD Configuration Retrieval Method
Suppose that we are given a multilayer network following Def. 1. Our goal is to generate a digital twin, a multilayer network $\hat{G}$ generated by mABCD with associated graphs $\hat{G}_{i}=(\hat{V}_{i},\hat{E}_{i})$ for $i\in[\ell]$, that mimics the original network $G$ as best as possible. In order to do it, one needs to appropriately select the parameters $\Theta$ of mABCD and then measure the “success”, i.e., how close $G$ to $\hat{G}$ is. The success is measured by computing a divergence score $\mathcal{D}(\hat{G},G)$; the smaller the score, the better the fit is obtained.
@@
3.0.1 Dimension of reference latent layer ($d$).
Communities in mABCD are independently generated in each layer, but the desired correlations are obtained via a hidden, low-dimensional geometric reference layer. Nodes close to each other in the reference layer end up in the same community with a larger probability. In particular, the dimension of a reference layer $d$ affects many properties of the generated network (see Sec. 4 in [17]). However, it seems not so easy to guess the right value so we leave it as a parameter that eventually can be tuned for the best outcome, that is, the smallest divergence score $\mathcal{D}(\hat{G},G)$. Understanding the influence of this hyperparameter and selecting the optimal value is left as an open problem (Open Problem 2).
This geometric approach to community structure is justified by the fact that latent geometric spaces are widely believed to shape complex networks (e.g., social media networks shaped by users’ opinions, education, knowledge, interests, etc.). These latent spaces have been successfully employed for many years to model and explain network properties such as self-similarity [21], homophily, and aversion [10]. For further references, we direct the reader to the survey [4] or the book [20].
@@
To measure correlations between edges in different layers, we define $\mathbf{R}$, an $\ell\times\ell$ matrix in which elements $\mathbf{R}_{i,j}\in[0,1]$ ($i,j\in[\ell]$) capture correlation between edges present in layers $i$ and $j$. For any $i,j\in[\ell]$, let $E_{i}^{j}$ be the set of edges that are present in layer $i$, involving actors that are also active in layer $j$. Entries in $\mathbf{R}$ are computed using the following formula:
$\mathbf{R}_{i,j}=\frac{|E_{i}^{j}\cap E_{j}^{i}|}{\min\{|E_{i}^{j}|,|E_{j}^{i}|\}}.$
If $\min\{|E_{i}^{j}|,|E_{j}^{i}|\}=0$, then we leave $\mathbf{R}_{i,j}$ undefined.
Note that $\mathbf{R}_{i,i}=1$ for any $i\in[\ell]$ and $\mathbf{R}_{i,j}=\mathbf{R}_{j,i}$ for $1\leq i<j\leq\ell$. The maximum value of $1$ is attained when edges in one of the layers form a subset of edges in the other layer. The minimum value of 0 is attained when the two sets of edges in the corresponding layers are completely disjoint.
Extracting $\mathbf{R}$ from the original network $G$ is straightforward. The mABCD model aims to produce a synthetic network with the desired edge correlation matrix, but this is not guaranteed, and some small error always occurs. Let $\hat{\mathbf{R}}$ be the corresponding edge correlation matrix for the obtained digital twin $\hat{G}$. The divergence score is simply the normalised to $[0,1]$ Frobenius norm between the two matrices:
$\mathcal{D}_{\mathbf{R}}(\hat{G},G)=\frac{\lVert{\mathbf{R}}-\hat{\mathbf{R}}\rVert_{F}}{\sqrt{\ell(\ell-1)}}=\sqrt{\frac{1}{\ell(\ell-1)}\sum_{i,j\in[\ell]}(\mathbf{R}_{ij}-\hat{\mathbf{R}}_{ij})^{2}}=\sqrt{\frac{1}{\binom{\ell}{2}}\sum_{1\leq i<j\leq\ell}(\mathbf{R}_{ij}-\hat{\mathbf{R}}_{ij})^{2}}.$
@@
To estimate $\tau_{i}$, which relies on the ordering of actor labels, we first sort actors with respect to their total degree (sum of degrees over all layers), breaking ties randomly if needed; that is, the actor with the highest total degree is assigned label $n$, the highest possible label. Then, for any $i\in[\ell]$, we compute the Kendall rank correlation coefficient $\tau_{i}$ between these labels and the corresponding degree sequence in layer $i$. Moreover, only nodes with a positive degree are taken into account to obtain the correlation value in order to reduce the noise; i.e., inactive nodes are ignored as they should be.
To measure the quality of the fit, one can compute an $\ell\times\ell$ matrix $\mathbf{A}$, where $\mathbf{A}_{ij}\in[-1,1]$ is the Kendall rank correlation coefficient between layers $i$ and $j$ for the real network $G$. Clearly, only nodes that are active in both layers (i.e., nodes in $V_{i}\cap V_{j}$) are considered. Once a digital twin is generated, matrix $\hat{\mathbf{A}}$ is computed analogously, but this time for the twin $\hat{G}$. As before, the divergence score is simply the (normalised) Frobenius norm between the two matrices:
$\mathcal{D}_{\tau}(\hat{G},G)=\frac{\lVert\mathbf{A}-\hat{\mathbf{A}}\rVert_{F}}{\sqrt{4\ell(\ell-1)}}=\sqrt{\frac{1}{4\binom{\ell}{2}}\sum_{1\leq i<j\leq\ell}(\mathbf{A}_{ij}-\hat{\mathbf{A}}_{ij})^{2}}.$
@@
This family of parameters, which models correlations between communities formed by nodes that are active in two different layers, appears to be the most challenging to retrieve. That results from inaccessibility of the latent biscuit-like reference layer (whence the communities are sampled and whose dimension is modelled by $d$) used during network generation. However, once a digital twin is generated, the success of estimating $r$ can be assessed by computing the corresponding divergence score $\mathcal{D}_{r}(\hat{G},G)$, which measures a feature indirectly captured by $r$ — the cross-layer alignment of communities between $G$ and $\hat{G}$. By efficiently searching the parameter space to identify the value $\hat{r}$ that minimises $\mathcal{D}_{r}(\hat{G},G)$, one can attempt to fit this parameter.
Let us then concentrate on the divergence score. For each layer $i\in[\ell]$ in the real network $G$, we independently run some stable clustering algorithm (in experiments, the Greedy Modularity Optimisation [6]) to get a partition $\mathcal{P}_{i}$ that identifies its community structure. Then, we compute an $\ell\times\ell$ matrix $\mathbf{B}$, where $\mathbf{B}_{ij}\in[0,1]$ is the adjusted mutual information (AMI) between partitions $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ induced by nodes that are active in both layers, that is, partitions induced by $V_{i}\cap V_{j}$. Matrix $\hat{\mathbf{B}}$ is computed analogously but for the twin $\hat{G}$ and the divergence score is defined as follows:
$\mathcal{D}_{r}(\hat{G},G)=\frac{\lVert\mathbf{B}-\hat{\mathbf{B}}\rVert_{F}}{\sqrt{\ell(\ell-1)}}=\sqrt{\frac{1}{\binom{\ell}{2}}\sum_{1\leq i<j\leq\ell}(\mathbf{B}_{ij}-\hat{\mathbf{B}}_{ij})^{2}}.$
The proposed method for predicting $r$ is based on access to the previously estimated parameters of the mABCD configuration ($\Theta$) and the $\mathbf{B}$ matrix of the original network. The objective function generates $n$ candidate twins from the configuration $\Theta^{(r)}$ corresponding to the evaluated value of $r$. It then computes a $\hat{\mathbf{B}}$ matrix for each twin and returns the divergence score $\mathcal{D}_{r}$ between $\mathbf{B}$ and the mean matrix obtained from the collection of $\hat{\mathbf{B}}$ matrices, denoted by $\bar{\mathbf{B}}$. This value is subsequently minimised during the Bayesian optimisation process to obtain the best-matching vector $\hat{r}$:
$\hat{r}=\arg\min_{r\in[0,1]^{\ell}}\mathcal{D}_{r}(\mathbf{B},\,\bar{\mathbf{B}}):\ \bar{\mathbf{B}}=\frac{1}{n}\sum_{k=1}^{n}\hat{\mathbf{B}}^{(k)}(\Theta^{(r)})$
By sequentially evaluating the objective function at values of $r$ selected based on previous evaluations, the method seeks to identify the best-matching value of $r$. The implementation employed, provided by the skopt [19] library, accounts for the non-deterministic behaviour of mABCD by incorporating Gaussian noise into the surrogate model of the objective. A typical optimisation trajectory for fitting $r$ is shown in Fig. 2.
Figure 2: A typical optimisation trajectory of $r$ projected on $\mathbb{R}^{2}$ space obtained during experiments. Note the preserved trade-off between exploration and exploitation, by which the employed method avoids getting stuck in local minima.
@@
The mABCD model generates, by design, power-law sequences with parameters provided as the input. Hence, the only discrepancy between the original network $G$ and its synthetic twin $\hat{G}$ comes from fitting the power-law distribution to the observed degree sequence of the real network. To measure the quality of the fit, we employ the Kolmogorov–Smirnov test. It is a nonparametric statistical test used to determine if a sample comes from a specific distribution (one-sample) or if two samples come from the same distribution (two-sample) by measuring the maximum difference between their cumulative distribution functions. This leads to the following divergence score:
@@
For simplicity, in this paper we assume that the degree sequences within layers of a real multilayered network $G$ (i.e., specific to nodes rather than actors) approximately follow a power law. But not all networks have degree sequences of this nature. To solve this issue, the mABCD model is flexible and allows the degree sequences to be directly injected into the model. Since mABCD is using the classical configuration model to generate edges in each layer, with this more advanced option, the degree sequence of the digital twin $\hat{G}$ would match exactly the original network $G$. We leave an implementation of this extension into our framework and an investigation of its consequences on the quality of the digital twin for the future (Open Problem 3).